# Enhancing Repository-Level Code Generation with Integrated Contextual Information

### v1

@@ Abstract.
Large language models (LLMs) have demonstrated remarkable capabilities in code generation tasks. However, repository-level code generation presents unique challenges, particularly due to the need to utilize information spread across multiple files within a repository. Existing retrieval-based approaches sometimes fall short as they are limited in obtaining a broader and deeper repository context. In this paper, we present CatCoder, a novel code generation framework designed for statically typed programming languages. CatCoder enhances repository-level code generation by integrating relevant code and type context. Specifically, it leverages static analyzers to extract type dependencies and merges this information with retrieved code to create comprehensive prompts for LLMs. To evaluate the effectiveness of CatCoder, we adapt and construct benchmarks that include 199 Java tasks and 90 Rust tasks. The results show that CatCoder outperforms the RepoCoder baseline by up to 17.35%, in terms of pass@$k$ score. Furthermore, the generalizability of CatCoder is assessed using various LLMs, including both code-specialized models and general-purpose models. Our findings indicate consistent performance improvements across all models, which underlines the practicality of CatCoder.
@@
copyright: none
Figure 1. A scenario of developing a Java method triu in the Apache Commons Math (com, 2024) library. (1) The left part illustrates relevant code, whose patterns might be helpful when developers write current method code. (2) The right part illustrates the type context (i.e., the fields and method signatures of a class) related to the current method, which can be derived iteratively through static analysis, starting with the method signature.
@@ 1. Introduction
Large language models (LLMs) (Wang et al., 2021; Nijkamp et al., 2022; Roziere et al., 2023; Lozhkov et al., 2024) have been widely applied in various software engineering-related fields, and they have shown remarkable performance in code generation tasks (Chen et al., 2021; Austin et al., 2021). Code generation tasks can be divided into two types: standalone code generation and repository-level code generation. For the standalone scenario, the function can be invoked independently, and the generation of LLMs is conditioned only on the in-file context (i.e., contents within the same file). However, for more practical repository-level code generation, LLMs struggle to generate correct code if only in-file context is provided, due to that relevant information usually lies in many different files in a repository, and the lack of cross-file context is likely to lead to hallucinations. Therefore, it is essential to fully utilize the repository information and provide cross-file contexts for the LLMs to generate the correct repository-level code.
@@
To utilize repository information, existing studies use retrieval-based approaches to find similar code for reference. Zan et al. (Zan et al., 2022) train a dual-encoder model to retrieve API information from library documentation. Lu et al. (Lu et al., 2022) build an external code database to retrieve similar code. Shrivastava et al.  (Shrivastava et al., 2023) propose RLPG, in which they train a classifier to predict prompt proposals taken from repository files. Zhang et al. (Zhang et al., 2023) propose RepoCoder, which iteratively retrieves similar code in a repository. However, these approaches still have limitations. First, for approaches that use external sources or train models, they are not generalizable to new unseen repositories. Second, retrieved code cannot always cover the necessary logic and API invocations that correct code requires, especially when a code repository does not contain much duplicated code (Zhang et al., 2023).
To address the latter limitation, we find that apart from code retrieval, in statically typed programming languages [^1] Statically typed programming languages are those where the data types of variables have to be declared during compile-time, such as C, C++, Java and Rust., the types related to a function can also provide valuable references to LLMs. In other words, they can provide a set of fields and methods for the LLM to select and use in the generated code. In this paper, we refer to this information as type context. The type context serves as a complement to the code retrieval and helps LLMs better understand the repository structure at a deep level. To extract such information, we utilize existing static analyzers.
@@
Based on type context, we propose a new repository-level code generation framework named CatCoder, which exploits the combination of relevant code and type context for statically typed languages. It works on top of a frozen LLM and does not require external databases or model training. First, it retrieves relevant code from the repository. Then, CatCoder invokes a static analyzer to obtain dependent types for the function and creates textual representations of the type context. Finally, CatCoder merges two kinds of information into the prompt and generates the function code with an LLM.
@@
To evaluate the effectiveness of CatCoder, we adapt and construct repository-level code generation benchmarks that include 199 Java tasks and 90 Rust tasks in total. We compare CatCoder with vanilla LLM, the In-File context, and RepoCoder (Zhang et al., 2023), a state-of-the-art framework for repository-level code generation, in terms of pass@$k$ (Chen et al., 2021) scores. To validate the generalizability of CatCoder, we evaluate CatCoder with various LLMs, including code-specialized models and general purpose models. The results on the benchmarks indicate that CatCoder substantially outperforms the vanilla LLM and the In-File context. In addition, it outperforms RepoCoder by up to 17.35%. The results of the generalizability study demonstrate that CatCoder improves the performance of all selected LLMs.
In summary, we make the following contributions:
@@
Paper Organization: The remainder of the paper is structured as follows. Section 2 introduces the motivation of CatCoder using an example. Section 3 presents the details of CatCoder. Section 4 describes the experimental setups, including baselines, evaluation metric, benchmarks, and implementation details. Section 5 analyzes the experimental results and provides answers to the research questions. Section 6 discusses the failure cases of CatCoder and threats to validity. Section 7 summarizes the related work. Section 8 concludes the paper with possible future work.
Figure 2. Overview of our approach
@@ 2. Motivating Example
In this section, we present a motivating example to illustrate the main idea of CatCoder. Figure 1 shows a scenario for developing a Java method named triu|, which takes a RealMatrix object and an int as input and returns a RealMatrix object.
From the perspective of human developers, to write repository-level code, we do not only refer to the method signature and the docstrings but also search for relevant code as references. In addition, for statically typed programming languages like Java, it is common to dive into related types (e.g., argument and return types) to seek available APIs. The two kinds of information can be concluded as follows:
@@
    Relevant code with a similar name, parameters, return type, or docstrings. As similar code is more likely to perform similar or symmetric operations, we can learn from the patterns and write new code.
@@
    Type context, i.e., the field and method information of related classes, which provides hints on the possible ways to assign values or invoke APIs.
[…]
After obtaining such information, developers can write the code by selecting useful parts of the relevant code and the type context and combining them with some logic. Given the powerful in-context learning ability of LLMs, it is intuitive that LLMs can also benefit from such comprehensive information if we utilize it to augment the prompts.
To improve the performance of repository-level code generation with LLMs, existing work (Lu et al., 2022; Zhang et al., 2023) uses retrieval-augmented generation to retrieve relevant code as presented in Info.1. However, it is limited as it ignores Info.2 (i.e., type context). For code repositories that do not contain many similar method pairs, the retrieved code would be less helpful. In this case, without type context, LLMs are more likely to hallucinate and generate non-existent field accesses or API calls.
Motivated by the aforementioned example, we illustrate the idea of our work CatCoder, which highlights the combination of Info.1 and Info.2. CatCoder is carried out in the following steps:
@@ 3. Proposed Approach
The framework of CatCoder is presented in Figure 2. CatCoder takes a source code repository and an empty function (along with its docstrings, if it exists). The output of CatCoder is the completed code of the function.
In addition to Figure 2, Algorithm 1 presents the pseudocode of the workflow of CatCoder. CatCoder contains two core components: (1) Relevant Code Retrieval and (2) Type Context Extraction. By utilizing these components, CatCoder is able to gather two kinds of referential information for the LLM. After post-processing the raw output of the LLM, CatCoder returns the generated code.
@@ 3.1.1. Code Splitting
We traverse all source code files in the repository and split them into smaller chunks. Starting with complete files, we set a maximum chunk size and recursively split larger fragments into smaller ones that fit in the chunk.
As the structure of the source code is often complex, we cannot split it at an arbitrary position. Otherwise, the resulting contents would be less meaningful and difficult to understand. Therefore, to better preserve the code structures, we split the code only before the following positions (sorted by priority in descending order):
@@
    Type definitions
@@
    Function definitions
@@
    Control flow statements
@@
    New lines
After splitting, we obtain a corpus of code chunks [^2] We call them “documents” later in this subsection. to perform retrieval.
@@ 3.1.2. Hybrid Retrieving
[…]
@@
We choose bag-of-words model for sparse retrieval, and use the BM25 function (Robertson et al., 1995; Trotman et al., 2012) to rank the set of documents, based on the query terms. Given a document $D$ and a query $Q$ with terms $t_{1},t_{2},\cdots,t_{n}$, the BM25 score can be defined as follows:
[…]
where $\mathrm{IDF}(t,D)$ is the inverse document frequency weight, $\mathrm{TF}(t,D)$ is the term frequency (i.e., the number of appearing times in the document), $|D|$ is the length of the document, and $k_{1}$ and $b$ are hyperparameters. In default, the values of $k_{1}$ and $b$ are set to 1.5 and 0.75. $\mathrm{IDF}(t,D)$ can be calculated by:
$\mathrm{IDF}(t,D)=\log\dfrac{N_{D}-\mathrm{DF}(t,D)+0.5}{\mathrm{DF}(t,D)+0.5},$
where $N_{D}$ is the number of documents and $\mathrm{DF}(t,D)$ is the document frequency (i.e., the number of documents that contain the term).
In practice, after splitting the code files from a repository, we apply the aforementioned BM25 function for a query and obtain a list of scores. After that, we select documents with top-$k$ highest scores.
@@
$d^{2}(\mathbf{p},\mathbf{q})=\left\|\mathbf{p}-\mathbf{q}\right\|^{2}=\sum_{i=1}^{n}(p_{i}-q_{i})^{2},$
where $n$ is the number of dimensions in the embedding vectors. A lower value of $d^{2}(\mathbf{p},\mathbf{q})$ indicates that $\mathbf{p}$ and $\mathbf{q}$ are more similar.
@@ 3.1.3. Result Ranking
In the third step, we ensemble the two retrievers and re-rank their outputs using the Reciprocal Rank Fusion (RRF) (Cormack et al., 2009) algorithm. We assign weights to the retrievers and calculate the weighted RRF score [^4] We assume that a retriever does not output duplicate documents. However, duplicate documents across retrievers are acceptable. for each retrieved document $D$:
$\mathrm{RRF}(D)=\sum_{i=1}^{n}\mathbf{1}(D)\cdot\dfrac{w_{i}}{r(D)+60},$
where $w_{i}$ is the weight assigned to a certain retriever, $\mathbf{1}(D)$ is an indicator function [^5] The value is 1 if $D$ is one of the outputs of this retriever, otherwise the value is 0., and $r(D)$ is the original rank (if exists) of $D$ in the outputs of this retriever.
@@ 3.2. Type Context Extraction
Here, we present the type context and the way to extract type contexts from code repositories using static analyzers [^6] The word “type” is an abstract concept and might be different in various languages. For example, it refers to class, interface and enum in Java; whereas in Rust, it refers to struct and enum..

### v2

# CatCoder: Repository-Level Code Generation with Relevant Code and Type Context

@@ Abstract.
Large language models (LLMs) have demonstrated remarkable capabilities in code generation tasks. However, repository-level code generation presents unique challenges, particularly due to the need to utilize information spread across multiple files within a repository. Specifically, successful generation depends on a solid grasp of both general, context-agnostic knowledge and specific, context-dependent knowledge. While LLMs are widely used for the context-agnostic aspect, existing retrieval-based approaches sometimes fall short as they are limited in obtaining a broader and deeper repository context. In this paper, we present CatCoder, a novel code generation framework designed for statically typed programming languages. CatCoder enhances repository-level code generation by integrating relevant code and type context. Specifically, it leverages static analyzers to extract type dependencies and merges this information with retrieved code to create comprehensive prompts for LLMs. To evaluate the effectiveness of CatCoder, we adapt and construct benchmarks that include 199 Java tasks and 90 Rust tasks. The results show that CatCoder outperforms the RepoCoder baseline by up to 14.44% and 17.35%, in terms of compile@$k$ and pass@$k$ scores. In addition, the generalizability of CatCoder is assessed using various LLMs, including both code-specialized models and general-purpose models. Our findings indicate consistent performance improvements across all models, which underlines the practicality of CatCoder. Furthermore, we evaluate the time consumption of CatCoder in a large open source repository, and the results demonstrate the scalability of CatCoder.
@@
$c$s: Software and its engineering Automatic programming
@@ 1. Introduction
Large language models (LLMs) (wang2021codet5; nijkamp2022codegen; codellama; starcoder2) have been widely applied in various software engineering-related fields, and they have shown remarkable performance in code generation tasks (humaneval; mbpp). Recently, repository-level code generation has attracted much attention (liu2023repobench; wu2024repoformer; zhang2023repocoder; shrivastava2023repository). It refers to the generation of a specific function within a repository and is thus more practical than standalone code generation tasks (i.e., generating functions with no dependencies). This practical scenario introduces unique challenges, as the target code often depends on definitions and usage patterns distributed across multiple files, rather than being confined to a single file. When provided with only the in-file context for such a task, an LLM may hallucinate and produce code that appears plausible but fails to compile or execute correctly within the project (e.g., misusing an API or referencing an undefined identifier). Therefore, a key problem in repository-level code generation is how to provide the LLM with the necessary cross-file context so that it can generate valid and consistent code.
@@
Existing studies have tried to tackle this problem by utilizing repository information with retrieval-based approaches. Zan et al. (private_library) train a dual-encoder model to retrieve API information from library documentation. Lu et al. (lu2022reacc) build an external code database to retrieve similar code. Shrivastava et al.  (shrivastava2023repository) propose RLPG, in which they train a classifier to predict prompt proposals taken from repository files. Zhang et al. (zhang2023repocoder) propose RepoCoder, which iteratively retrieves similar code in a repository. However, these approaches still have limitations. First, methods that depend on external resources or require training specialized retrieval models often lack generalizability across unseen repositories or programming languages. Second, the retrieved code may fail to capture all necessary field accesses and API calls required for correct generation. This issue is particularly pronounced in repositories with minimal code duplication (zhang2023repocoder), or when the desired functionality is novel within the codebase. These limitations underscore the need for more comprehensive modeling of repository context.
In this paper, we aim to address the aforementioned limitations by capturing a broader repository context for the LLM. We observe that, in addition to code retrieval, the types associated with a function can also serve as valuable references for LLMs. Specifically, these related types provide a rich set of fields and methods from which the LLM can select when generating code, thereby enhancing its understanding of repository-specific details. We refer to this explicit information of types as type context. The type context encapsulates localized knowledge embedded in nested scopes, serves as a complement to code retrieval, and helps LLMs develop a deeper understanding of the repository structure. Our work focuses on constructing type context for statically typed programming languages [^1] Statically typed programming languages require variable types to be declared at compile time, such as C, C++, Java, and Rust., where precise type information can be reliably extracted and leveraged. In contrast, dynamically typed languages (e.g., Python and JavaScript) determine variable types at runtime, making type information less accessible and more ambiguous. To extract type context, we utilize existing static analyzers.
@@
Building on the concept of type context, we propose a new repository-level code generation framework named CatCoder. By leveraging the combination of relevant code and type context, CatCoder aims to improve the generation of code that is both logically sound and correctly integrated within its local environment. The framework operates in three stages. First, it retrieves relevant code from the repository. Next, it invokes a static analyzer to obtain dependent types for the target function and constructs a textual representation of the type context. Finally, it integrates both the retrieved code and the type context into a single prompt and generates the target function using an LLM. CatCoder runs on top of a frozen LLM and does not require external databases or additional model training.
@@
To evaluate the effectiveness of CatCoder, we adapt and construct repository-level code generation benchmarks that include 199 Java tasks and 90 Rust tasks in total. We compare CatCoder with vanilla LLM, the In-File context (i.e., contents within the same file), and RepoCoder (zhang2023repocoder), a state-of-the-art framework for repository-level code generation, in terms of compile@$k$ and pass@$k$ (humaneval) scores (i.e., compilation rate and test passing rate). To validate the generalizability of CatCoder, we evaluate CatCoder with various LLMs, including code-specialized models and general-purpose models. To assess the scalability of CatCoder, we conduct extensive evaluations on a large open-source repository, measuring the time consumption of sequential code generation tasks. The results on the benchmarks indicate that CatCoder substantially outperforms the vanilla LLM and the In-File context. In addition, it outperforms RepoCoder by up to 14.44% and 17.35%, in terms of compile@$k$ and pass@$k$. The results of the generalizability study demonstrate that CatCoder improves the performance of all selected LLMs. Furthermore, the results of the scalability study underline the efficiency and practicality of CatCoder in large repositories.
Contributions. In summary, we make the following contributions:
@@
  4. (4)
    We release the replication package [^2] https://github.com/pan2013e/catcoder of CatCoder, including the source code of CatCoder and our benchmarks, to facilitate future research.
Paper Organization. The remainder of the paper is structured as follows. Section 2 introduces the motivation of CatCoder using an example. Section 3 presents the details of CatCoder. Section 4 describes the experimental setups, including baselines, evaluation metrics, benchmarks, and implementation details. Section 5 analyzes the experimental results and provides answers to the research questions. Section 6 discusses the failure cases of CatCoder and threats to validity. Section 7 summarizes the related work. Section 8 concludes the paper with possible future work.
Figure 1. The two essential sources of local context for generating the triu method. (1) Relevant Code (the ones method) provides a crucial, repository-specific pattern for instantiating a RealMatrix object. (2) Type Context provides the equally crucial API definition for the RealMatrix interface, including methods like getEntry needed for element access.
@@ 2. Motivating Example
In this section, we present a realistic development scenario to illustrate the core idea behind CatCoder. The motivating example will demonstrate that synthesizing information from both relevant code and type context is crucial for successful generation.
Consider the task of implementing the helper method triu within the CMAESOptimizer class from the widely used Apache Commons Math library, as shown in Fig. 1. As described in its docstring, the triu method is intended to extract the upper triangular part of a given matrix. Its signature is private static RealMatrix triu(final RealMatrix m, int k), where $m$ is the input matrix and $k$ specifies the diagonal offset.
Based on its context-agnostic knowledge of coding, an LLM can infer the general algorithmic principle: the implementation should return a new matrix with the same dimensions, where an element at position $(r,c)$ is copied from $m$ if the column index $c$ is greater than or equal to $r+k$; otherwise, the element is set to zero. However, the requirement of context-dependent knowledge poses challenges to the LLM. The RealMatrix type is an interface in Apache Commons Math and cannot be treated as a nested double array. Proper interaction with a RealMatrix instance requires familiarity with repository-specific API conventions, such as using getEntry(row, column) to access elements. Furthermore, creating a new RealMatrix instance requires using a specific concrete implementation, Array2DRowRealMatrix, and understanding the correct constructor usage. An LLM must therefore grasp these repository-specific details to produce valid code.
Table 1. Different information sources for the triu task
Information Source Key Insight Provided Sufficient Alone?
Info.1 ones method Idiomatic instantiation:
new Array2DRowRealMatrix(...)
gray[1pt/1pt]blackInfo.2 RealMatrix API Correct API access:
m.getRow/ColumnDimension(...)
m.getEntry(...)
gray[1pt/1pt]blackCatCoder (Info.1 + Info.2) Complete, correct implementation
To generate a correct implementation, an LLM must solve two distinct subproblems, that is, (1) how to create the new RealMatrix to be returned, and (2) how to read data from the input RealMatrix m:
  1. (1)
    As shown in Fig. 1, the ones method is retrieved due to its shared return type and location. From this snippet, an LLM can learn a non-trivial pattern for object creation: instantiate a new Array2DRowRealMatrix by first creating a primitive double array and passing it to the constructor. This form of relevant code retrieval reflects a common practice among developers, who often refer to similar methods when writing repository-level code.
  2. (2)
    In addition, Fig. 1 illustrates the available APIs of the RealMatrix type, presented as the type context. Because both the target method triu and the surrounding class CMAESOptimizer reference RealMatrix, its API information provides essential insights. For instance, an LLM can learn how to access dimensions and elements correctly by referencing the provided method signatures. This mirrors another common development strategy: navigating related types (e.g., parameter or return types) to discover usable fields and methods.
Table 2 categorizes the two aforementioned information sources and their key insights. We abstract them as follows:
@@
    Relevant code snippets that share similarities in method name, parameters, return types, or docstrings. These examples often perform semantically related or symmetric operations and thus provide reusable implementation patterns.
@@
    Type context, defined as the available fields and methods of related types, offering clues for accessing or manipulating object instances correctly.
This motivating example highlights that both Info.1 and Info.2 are indispensable. The absence of either is likely to lead to incorrect or incomplete code generation. While prior work (lu2022reacc; zhang2023repocoder) has leveraged retrieval-augmented generation to utilize relevant code (Info.1), they typically overlook Info.2 (type context). As Fig. 1 shows, retrieved code alone may not fully capture the required API usage. Moreover, in repositories that lack code duplication, relying solely on retrieval is less effective. In such cases, the absence of type context increases the risk of hallucination, such as referencing non-existent methods or fields. To address this gap, CatCoder integrates both Info.1 and Info.2. Our framework is carried out in the following steps:
@@ 3. Proposed Approach
The framework of CatCoder is presented in Fig. 2. CatCoder takes a source code repository and an empty function (along with its docstrings, if it exists). The output of CatCoder is the completed code of the function.
In addition to Fig. 2, Algorithm 1 presents the pseudocode of the workflow of CatCoder. CatCoder contains two core components: (1) Relevant Code Retrieval and (2) Type Context Extraction. By utilizing these components, CatCoder is able to gather two kinds of referential information for the LLM. After post-processing the raw output of the LLM, CatCoder returns the generated code.
@@ 3.1.1. Code Splitting
We traverse all the source code files in the repository and split them into smaller chunks. Starting with complete files, we set a maximum chunk size and recursively split larger fragments into smaller ones that fit in the chunk.
As the structure of the source code is often complex, we cannot split it at an arbitrary position. For example, we should not break in the middle of a statement line or some important code elements (e.g., function definitions). Otherwise, the resulting contents would be less meaningful and difficult to understand. Therefore, we split the code only before the following positions (sorted by priority in descending order):
@@
    Type definitions (e.g., public class Main in Java)
@@
    Function definitions (e.g., public static void main(String[] args) in Java)
@@
    Control flow statements (e.g, if, for and while statements in Java)
@@
[…]
This rule follows the hierarchy of many programming languages and helps protect code structures during splitting. After splitting, we obtain a corpus of code chunks [^3] We call them “documents” later in this subsection. to perform retrieval.
@@ 3.1.2. Hybrid Retrieving
In the second step, we build retrievers to search for relevant code. To achieve better search performance, we follow previous research (luan2021sparse) and ensemble two kinds of retrievers: a sparse one and a dense one. The sparse retriever corresponds to keyword matching, and the dense retriever corresponds to semantic search. Each retriever searches and returns top-kk444kk is a hyperparameter, and we describe the settings in Section 4.4.4. relevant documents (i.e., split code chunks), after a query is constructed by concatenating docstrings and the function signature.
@@
We choose the bag-of-words model for sparse retrieval, and use the BM25 function (robertson1995okapi; trotman2012towards) to rank the set of documents, based on the query terms. Given a document $D$ and a query $Q$ with terms $t_{1},t_{2},\cdots,t_{n}$, the BM25 score can be defined as follows:
BM25(D, Q) = ∑_t ∈Q IDF(t,D) ⋅TF(t,D) ⋅(k1+1)TF(t,D) + k1⋅(1-b+b⋅—D——D—avg) where $\mathrm{IDF}(t,D)$ is the inverse document frequency weight, $\mathrm{TF}(t,D)$ is the term frequency (i.e., the number of appearing times in the document), $|D|$ is the length of the document, and $k_{1}$ and $b$ are hyperparameters. By default, the values of $k_{1}$ and $b$ are set to 1.5 and 0.75. $\mathrm{IDF}(t,D)$ can be calculated by:
IDF(t, D) = logND-DF(t, D) + 0.5DF(t, D) + 0.5, where $N_{D}$ is the number of documents and $\mathrm{DF}(t,D)$ is the document frequency (i.e., the number of documents that contain the term).
In practice, after splitting the code files from a repository, we apply the aforementioned BM25 function for a query and obtain a list of scores. After that, we select documents with the top-$k$ highest scores.
@@
d^2(p,q) = ∥p - q ∥^2 = ∑_i=1^n (p_i - q_i)^2, where $n$ is the number of dimensions in the embedding vectors. A lower value of $d^{2}(\mathbf{p},\mathbf{q})$ indicates that $\mathbf{p}$ and $\mathbf{q}$ are more similar.
@@ 3.1.3. Result Ranking
In the third step, we ensemble the two retrievers and re-rank their outputs using the Reciprocal Rank Fusion (RRF) (rrf) algorithm. We assign weights to the retrievers and calculate the weighted RRF score [^5] We assume that a retriever does not output duplicate documents. However, duplicate documents across retrievers are acceptable. for each retrieved document $D$:
RRF(D) = ∑_i=1^n 1(D)⋅wir(D) + 60, where $w_{i}$ is the weight assigned to a certain retriever, $\mathbf{1}(D)$ is an indicator function [^6] The value is 1 if $D$ is one of the outputs of this retriever, otherwise the value is 0., and $r(D)$ is the original rank (if exists) of $D$ in the outputs of this retriever.
@@ 3.2. Type Context Extraction
Here, we present the type context and the way to extract type contexts from code repositories using static analyzers [^7] The word “type” is an abstract concept and might be different in various languages. For example, it refers to class, interface, and enum in Java; whereas in Rust, it refers to struct and enum..